OpenStudy (anonymous):

There are two distinct round tables, each with 5 seats. In how many ways may a group of ten people sit?

OpenStudy (anonymous):

i am going to go out on a limb and say \[\dbinom{10}{5}4!4!\] then you can tell me why this is wrong

OpenStudy (anonymous):

well i haven't figured out the right answer for this one yet lol. but the answer is 145 152

OpenStudy (anonymous):

oh well let me check is this is the same! GOT IT!

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

now that i am confident i am not a total moron when it comes to counting, would you like the reasoning?

OpenStudy (anonymous):

yes please :)

OpenStudy (anonymous):

i reasoned as follows: the number of ways to put 5 people at one table and 5 at another is the same as the number of ways to choose 5 from a set of 10 which by definition is \[\dbinom{10}{5}\]

OpenStudy (anonymous):

and \[\dbinom{10}{5}=\frac{10\times 9\times 8\times 7\times 6}{5\times 4\times 3\times 2}=2\times 9\times 2\times 7=252\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

then once you have them divided in to two groups, the number of ways to arrange 5 people in a circle is 4! and since there are two tables you have 4! for one and 4! for the other so "counting principle" says multiply all this together

OpenStudy (anonymous):

btw there is a reasonable explanation for why if you have n people in a circle there are (n-1)! ways to arrange them, whereas if you have n people in a line there are n! ways

OpenStudy (anonymous):

and there is even a nice picture. if you want you can read the first paragraph here, which says it a lot better than i can http://www.learner.org/courses/mathilluminated/units/2/textbook/05.php first three pictures really says it all

OpenStudy (anonymous):

oh wow:) thanks a lot for your help